Using methyl iodide, CH3I, and sodium acetylide as the sources of carbon atoms, show how to synthesize meso-2,3-dibromobutane. You may use any other needed reagents or solvents, provided they do not supply carbon atoms to the final product. About 5 steps are required. Please write out the formulas of the reagents.Free Response (10 pts each) Read instructions carefully and write out your answers in detail. 1. Using methyl iodide, CHI, an
Answer #1

STEPS - + H-CEC Na - H-(=C-CH3 CH3 - NaNH2 chy- (=(-13 ang sa taong CH3-((-CH3 | w СЕС-СН3 H2 Pd | Basot Vind lae Catalyst 41